Logiweb dictionary of finalversion

Up Help

0 0 finalversion
1 0 pred calc
2 0 pc1
3 0 pc2
4 0 pc3
5 0 pc4
6 0 pc5
7 0 pc6
8 0 pc7
9 0 pc8
10 0 pc9
11 0 pc10
12 0 pc11
13 0 pc12
14 0 pcmp
15 0 pcunsound
16 0 pcia
17 0 pcie
18 0 pcdeduction
19 0 trivia
20 0 repeat
21 0 andintro
22 0 andelim1
23 0 andelim2
24 0 orintro1
25 0 orintro2
26 0 orelim
27 0 notintro
28 0 implyintro

29 0 notnotintro
30 0 notnotelim
31 0 mt
32 0 pbc
33 0 lem
34 0 forallintro
35 0 forallelim
36 0 existsintro
37 0 existselim
38 0 bottomelim
39 0 lemnotintro
40 0 hlplem1
41 0 hlplem2
42 0 hlplem3
43 0 hlplem4
44 0 hlplem5
45 0 goal1
46 0 hlplem6
47 0 goal2

48 2 ∗ ≡ ∗ 49 2 ∗ = ∗ 50 1 ¬∗ 51 2 ∗ ∧ ∗ 52 2 ∗ ∨ ∗ 53 2 ∀ ∗ . (∗) 54 2 ∃ ∗ . (∗) 55 2 ∗ ∈ ∗

The pyk compiler, version 0.grue.20060417+ by Klaus Grue GRD-2006-07-14.UTC:09:09:53.060274 = MJD-53930.TAI:09:10:26.060274 = LGT-4659585026060274e-6